Consider the following two investment alternatives: First, a risky portfolio that pays a 20% rate of return with a probability of 60% or a 5% rate of return with a probability of 40%. Second, a Treasury bill that pays 6%. If you invest $50,000 in the risky portfolio, your expected profit after one year would be
Answer #1

Risky Portfolio:

Return of 20% with a probability of 60% and a return of 5 % with a probability of 40%.

Expected Return of Risky Portfolio = Rp = 20 x 0.6 + 5 x 0.4 = 14 %

Investment Value = $ 50000

Expected Investment Value After One Year = 50000 x 1.14 = $ 57000

Expected Profit = 57000 - 50000 = $ 7000